A thermal tap used in certain apparatus consists of a silica rod which fits tightly
inside an Aluminium tube whose internal diameter is 8 mm at 0℃ . When the
temperature is raised, the fit is no longer exact. Calculate what change in temperature
is necessary to produce a channel whose cross-section is equal to that of a tube of
1 mm internal diameter
1
Expert's answer
2020-12-10T11:09:26-0500

Let's T1 = 273K, d1 = 8mm, d2 = 1mm;

we use the ideal gas law

PV/T = const

P1V1 / T1 = P2 V2 / T2

pressure is constant so

V1 / T1 = V2 / T2

"\pi"d12 / 4T1 = "\pi"d22 / 4T2

64 / T1 = 1 / T2

T2 = T1 / 64 = 273 / 64 = 4.23K;

the changing of temperature is that T2 - T1 = 268.67K;

the temperature decreasing to 268.67K